(1 pt) You wish to invest $1300 over one year in two accounts paying 5% and 6% annually. How much should you invest in each to earn $70?
Answer by richwmiller(17219)

x+y=1300,
0.05*x+0.06*y=70
x=1300-y
0.05*(1300-y)+0.06*y=70
65-0.05y+0.06*y=70
0.01*y=5
y= 500
x=1300-y
x=800
check
0.05*800+0.06*500=70
40+30=70
70=70
ok